OpenStudy (anonymous):

Suppose you have these two expressions: 1/2 = C -A -1/2 = A - C What is A =?, What is C =?

OpenStudy (anonymous):

When I try to equate the above, I get 0 = 0....

OpenStudy (anonymous):

lol fix it... say: suppose you have THIS expression: ... why? note. 1/2=C-A and -1/2=A-C is actually the same ... when we multiply both sides by -1 in the "second" equation we get 1/2=C-A

OpenStudy (anonymous):

Yes I understand that, but what would C, A be then, if the equations are exactly the same?

OpenStudy (anonymous):

so baisically subsitute one valu for the other right?

OpenStudy (anonymous):

you should Have at least 2 system of equations to find two different variabiles.

OpenStudy (anonymous):

if you'd have \[\LARGE \frac12 =C-A\] and \[\LARGE \frac34=4A-2C\] maybe we could do something... but without having two diffenrent equations. You can't.

OpenStudy (anonymous):

Both equations are exactly the same.

OpenStudy (anonymous):

yeah I know. That's the problem :(

OpenStudy (amorfide):

 no solution